OpenStudy (anonymous):

(BEST ANSWER GETS MEDAL!) A man started walking at 3 mph towards a city 72 miles away. Later, he was picked up by a motorist who took him to the city at 42 mph. If the entire trip took 4.5 hours, how far did the man walk? (JUST SHOW THE EQUATION YOU WOULD USE, NOT THE ANSWER)

OpenStudy (anonymous):

we can do this did you try it with numbers?

OpenStudy (anonymous):

no, I just need the equation for solving it

OpenStudy (anonymous):

Sorry, i just think it would be better for me to try to solve it myself :)

OpenStudy (anonymous):

lol that is the hard part, finding an equation once you have it solving is routine all the work comes in finding the equation i would start out with two variables, say x is the distance walked and y is the distance in the car, then you know \(x+y=72\) and you can also get an equation involving time, as it is distance divided by rate i can show you if you like or let you work on your own

OpenStudy (anonymous):

Can you show me? I'm suddenly quite confused xD @satellite73

OpenStudy (anonymous):

ok sure

OpenStudy (anonymous):

Yay thank you!

OpenStudy (anonymous):

we know the total is 72 right? so if we call the distance walked \(x\), which is what we have to solve for, we can either write \(x+y=72\) or we can just say the distance in the car is \(72-x\) either way

OpenStudy (anonymous):

Yes, that makes sense

OpenStudy (anonymous):

good now we have to deal with the time business we know time is distance divided by rate so the time walking is \(\frac{x}{3}\) and the time in the car is \(\frac{72-x}{42}\)

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

and to total time is \(4.5\) so \[\frac{x}{3}+\frac{72-x}{42}=4.5\] and we can solve for \(x\)

OpenStudy (anonymous):

Ooh that makes sense xD can I try to solve this and you tell me if I made any mistakes?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i give you a hint if you get a whole number, it is right

OpenStudy (anonymous):

I got x = 9, is that correct?

OpenStudy (anonymous):

@satellite73 also, I'm unsure of the way I solved it, how could show what you would do? :)

OpenStudy (anonymous):

yes it is 9

OpenStudy (anonymous):

Wait never mind, I understand how to do it!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

so whatever you did, it is right me, i would multiply both sides by \(42\) as step one to clear the fractions

OpenStudy (anonymous):

yep that's what I did! Thank you so much for your help!!!!

OpenStudy (anonymous):

yw !